# On the virialization threshold for halo mass functions

### v1

@@ Abstract
In a recent study by Euclid collaboration, the halo mass function (HMF) has been fitted with accuracy better than $1\%$ for the $\Lambda$CDM model. Several parameters were introduced and fitted against N-body simulations, assuming the usual linearly extrapolated matter density contrast at the collapse time, $\delta_{c}$, as a basic threshold for halo formation. As a result, a new function that multiplies $\delta_{c}$ was introduced, producing an effective threshold that varies both with redshift and mass scale. We show that the redshift evolution of this effective threshold is similar to the one of the linear extrapolated matter density contrast at the virialization time, $\delta_{\rm v}$. Assuming the Euclid HMF as a fiducial model, we refit the Sheth-Tormen (ST) HMF using $\delta_{\rm v}$ as a threshold. This new fit improves the agreement between ST-HMF and the Euclid one with respect to Despali et al. (2016) fit, specially at high masses. Interestingly, the parameters $a$ and $p$ in this refit have values closer to the Press-Schechter limit of the ST-HMF, showing that the use of $\delta_{\rm v}$ can provide semi-analytical HMF less dependent on extra parameters. Moreover, we analyze the consistency of the ST-HMF fitted with $\delta_{\rm v}$ in smooth dark energy models with time-varying equation of state, finding an overall good agreement with the evolution of halo abundances expected from the linear evolution of perturbations and the Euclid HMF extrapolated to these scenarios. These findings suggest that the use $\delta_{\rm v}$ as a basic function to describe the threshold for halo formation can be a good guide when considering extrapolations for models beyond $\Lambda$CDM, which are typically harder to study in simulations. We also provide a fitting formula for $\delta_{\rm v}$ in the $\Lambda$CDM model and a code to compute it for smooth dark energy described by the $w_{0}w_{a}$ parametrization.
@@ 1 Introduction
Since the first analytical proposal of a Halo Mass Function (HMF) by Press and Schechter (PS) [1], several semi-analytical and numerical studies on the topic have proposed alternative formulations, e.g., [2, 3, 4, 5, 6, 7, 8]. See [9] for an overview of several of these proposals. One of the key ingredients in the PS formulation, and in many others based in the Sheth-Tormen (ST) template [2], is the value of the critical density for the formation of halos, $\delta_{c}$, which mainly controls the exponential decay of the number of massive halos. Ideally, this parameter represents the value of the extrapolated linearly evolved matter contrast, $\delta_{m}$, above which halos are formed. For many years following the PS proposal, several distinct values for this threshold were used, e.g., [10, 11, 12]. Following the suggestions in [13, 14], the use of $\delta_{c}\simeq 1.686$, the value calculated for Einstein-de-Sitter (EdS) cosmology using the framework of the Spherical Collapse (SC) model [15, 16] at the moment of collapse, became standard.
@@
As suggest in [23, 24, 25], one can define another theoretical threshold for halo formation, $\delta_{{\rm v}}$, associated with virialization time. The definitions for the collapse and virialization threshold are:
@@
where $\delta_{m}^{L}$ is the linearly evolved matter perturbations at the redshift of collapse, $z_{c}$, or virialization, $z_{{\rm v}}$. Both of these quantities are computed in the framework of the SC model. For EdS, there is an analytical solution giving a constant $\delta_{{\rm v}}\simeq 1.583$. If one substitutes $\delta_{c}$ by $\delta_{{\rm v}}$ in HMFs, the threshold is reduced, as effectively done by the values of $a$ in the ST formulation, indicating the virialization time can provide a better description of the abundance of halos, especially at the high-mass tail of HMFs.
@@
In this work, we will further explore the possibility of using $\delta_{{\rm v}}$ instead of $\delta_{c}$ in HMFs, focusing in the $\Lambda$CDM model. First, we will show that the redshift-dependent part of the E-HMF is well described by a simple power-law function $\delta_{{\rm v}}$, with precision better $0.7\%$ for $0.2\leq\Omega_{m}^{0}\leq 0.4$, indicating that the correspondence between the functions is not a coincidence for a particular value of $\Omega_{m}^{0}$. Then we proceed with a reparametrization of the ST-HMF using $\delta_{{\rm v}}$ and assuming the E-HMF as a fiducial model, showing that this refit agrees better with E-HMF than the HMF fit given by [6]. As a result, we find that the refitted ST parameters $a$ and $p$ get closer to the values that reproduce the PS-HMF. We also analyse how the ST-HMF refited with $\delta_{{\rm v}}$ behaves when considering smooth DE models with $w\neq-1$.
The plan of this paper is as follows. In section 2 we review and discuss the collapse and virialization definitions for the thresholds. In section 3, we compare the redshift-dependent part of the E-HMF with $\delta_{{\rm v}}$ and reexpress $\Omega_{m}^{a_{z}}\delta_{c}^{2}$ as a function of $\delta_{{\rm v}}$. In section 4, substituting $\delta_{c}$ by $\delta_{{\rm v}}$, we refit the ST-HMF, assuming the E-HMF as a fiducial model. In section 5, we analyse the consistency of the ST-HMF refited with $\delta_{{\rm v}}$ for smooth DE models described by the $w_{0}w_{a}$ parametrization. We conclude in section 6 and provide a representation of $\delta_{{\rm v}}$ in terms of a well-known fit for $\delta_{c}$ for $\Lambda$CDM model in Appendix A, also indicating a repository where a code to compute $\delta_{{\rm v}}$ assuming the $w_{0}w_{a}$ parametrization can be found.
Throughout this paper, we assume as baseline the $\Lambda$CDM model with parameters: $h=0.7$, $\Omega_{c}^{0}h^{2}=0.125$, $\Omega_{b}^{0}h^{2}=0.022$, $A_{s}=2\times 10^{-9}$, $n_{s}=0.965$, $\tau=0.06$, $m_{\nu}=0.06$ and $\Omega_{k}^{0}=0$. In some examples, we vary $\Omega_{m}^{0}$ by changing only $\Omega_{c}^{0}h^{2}$. When analysing DE models described by the $w_{0}w_{a}$ parametrization, all the other parameters are kept fixed at the $\Lambda$CDM values. The computation of the matter power spectrum is made with CAMB [30].
@@ 2 Thresholds definitions
where $\rho_{m}\left(z\right)=\bar{\rho}_{m}\left(z\right)\left[1+\delta_{m}^{NL}\left(z\right)\right]$ is the total matter density. Note that $\Delta_{{\rm v}}$ is not the usual virialization overdensity, which is given by $\Delta=\rho_{m}\left(z_{\rm v}\right)/\bar{\rho}_{m}\left(z_{c}\right)\simeq 177.7\,.$
It is interesting to note that the virialization equation can be generalised to include other components, such as smooth or clustering DE [32, 33, 31, 26, 27], whereas the collapse definition is mostly insensitive to other ingredients that participate in the collapse process. Though in general $\delta_{m}^{NL}$ depends on the background evolution and other clustering components, near the collapse time, it is growing nearly exponentially, responding essentially to its own nonlinear value. Thus, the background evolution, and possibly other subdominant fluctuations, have very small impact on the value of $\delta_{c}$, typically less than $1\%$ with respect to the EdS value [28].
@@
When DE begins to dominate the expansion, they become redshift-dependent. Whereas $\delta_{c}$ has a slight decay when DE becomes important, $\delta_{{\rm v}}$ presents a mild growth, see figure 1.
@@
$\delta_{cM}=1.59+0.3114\times\ln\sigma_{8}\left(z\right)\,,$ (2.7)
@@
Figure 1: Evolution of $\delta_{c}^{2}$, $\delta_{cM}^{2}$, $\delta_{{\rm v}}^{2}$ and $a\delta_{c}^{2}$ as a function of redshift in the $\Lambda$CDM model with $\Omega_{m}^{0}=0.3$, $a_{z}=-0.0658$ and $a_{R}=0.7965$ (corresponding to a mass-scale $10^{14}M_{\odot}h^{-1}$ ).
@@ 3 The effective threshold in the E-HMF and $\delta_{{\rm v}}$
In this section, we analyse how the effective threshold of E-HMF is related to $\delta_{{\rm v}}$ and define the main quantities of interest. The comoving number density of halos per mass interval is given by
@@
where $\nu f\left(\nu\right)$ is the multiplicity function, $\nu=\delta_{c}/\sigma$, $\sigma^{2}$ is the smoothed variance of the matter power-spectrum, $P\left(k\right)$, given by
@@
$a=a_{R}\Omega_{m}^{a_{z}}=\left[a_{1}+a_{2}\left(\frac{d\ln\sigma}{d\ln R}+0.6125\right)^{2}\right]\Omega_{m}^{a_{z}}\,,$ (3.5)
$p=p_{1}+p_{2}\left(\frac{d\ln\sigma}{d\ln R}+0.5\right)\,,$ (3.6)
$q=\left[q_{1}+q_{2}\left(\frac{d\ln\sigma}{d\ln R}+0.5\right)\right]\Omega_{m}^{q_{z}}\,,$ (3.7)
@@
As can be seen, the function $a$ always multiplies $\delta_{c}^{2}$ and depends both on redshift and mass scale. The usual SC model is scale-independent, therefore it is not able to explain the variation with mass scale, which is encoded in $a_{R}$. Thus, we will focus on the redshift dependence of the effective threshold, $\Omega_{m}^{a_{z}}\delta_{c}^{2}$, and analyse its relation with $\delta_{{\rm v}}$.
@@
The evolution of $a\delta_{c}^{2}$ is shown in figure 1, where we used $a_{R}=0.7965$, corresponding to a mass of $10^{14}M_{\odot}h^{-1}$. We can see that both $\delta_{{\rm v}}^{2}$ and $a\delta_{c}^{2}$ increase at low-$z$, so they qualitatively represent the same trend for halo formation, i.e., a threshold that increases as $\Lambda$ becomes more important. Thus it is clear that the term $\Omega_{m}^{a_{z}}$ inside $a$ compensates the slight decay of $\delta_{c}$ to better describe the abundance of simulated halos. Moreover, $\delta_{{\rm v}}^{2}$ is closer to $a\delta_{c}^{2}$ then $\delta_{c}^{2}$. Therefore, also from a direct analysis of the E-HMF, one can conclude that the physics associated with the definition of $\delta_{{\rm v}}$ is more suitable than $\delta_{c}$ because it gives a value closer to the effective threshold and grows at low-$z$.
@@
However, it is also clear that there exists a mismatch between $\Omega_{m}^{a_{z}}\delta_{c}^{2}$ and $\delta_{\rm v}^{2}$. We can see that the former grows faster at late times and $a\delta_{c}^{2}$ has a lower value in the EdS limit. Our goal now is to find a representation of $\Omega_{m}^{a_{z}}\delta_{c}^{2}$ in terms of $\delta_{{\rm v}}$ and later check whether it is robust against variations of $\Omega_{m}^{0}$. We could compare $\Omega_{m}^{a_{z}}\delta_{c}^{2}$ and $\delta_{\rm v}^{2}$ directly, but as we will see this reparametrization will bring an important insight about the virialization model. Let us analyse the following description of the redshift-dependent part of the effective threshold
$\Omega_{m}^{a_{z}}\delta_{c}^{2}\left(z\right)=\alpha\delta_{{\rm v}}^{\beta}\left(z\right)\,,$ (3.8)
@@
where $\alpha$ and $\beta$ are constants. It is important to note that this form will introduce a non-standard power-law dependence on the density threshold, i.e., distinct from power $2$. This is unusual in the context of statistical derivation of HMF [1, 13], because the squared filtered density contrast has different power than its variance. One possible interpretation is that the extra powers of $\delta_{{\rm v}}$ encode the non-universality of HMF introduced by the new fitting functions. This distinct power can also be understood as a hint that the actual dependence of the threshold on the virialization process has stronger time-dependence than predicted by the SC model. As we will show, $\beta\simeq 4.37$ gives a representation of $\Omega_{m}^{a_{z}}\delta_{c}^{2}$ that is precise and robust against variations of $\Omega_{m}^{0}$. We checked that when keeping the power $2$ in the density threshold, the representation $\Omega_{m}^{a_{z}}\delta_{c}^{2}\left(z\right)=\alpha\Omega_{m}^{\beta}\delta_{{\rm v}}^{2}\left(z\right)$ gives better precision for a fixed $\Omega_{m}^{0}$, but is much less robust when varying $\Omega_{m}^{0}$.
@@
The constant $\alpha$ can be set by assuming that $\Omega_{m}^{a_{z}}\delta_{c}^{2}=\alpha\delta_{{\rm v}}^{\beta}$ in the EdS limit, thus
@@
Then $\beta$ is given by
$\beta=\ln\left(\frac{\Omega_{m}^{a_{z}}\delta_{c}^{2}\left(z\right)}{\delta_{c\,\text{EdS}}^{2}}\right)/\ln\left(\frac{\delta_{{\rm v}}\left(z\right)}{\delta_{{\rm v}\text{\,EdS}}}\right)\,.$ (3.10)
which varies with $z$. For $\Omega_{m}^{0}=0.3$ and $0\leq z\leq 2$, we have $4.514.05<\beta<$ 4.51.
@@
$\chi^{2}\left(\beta\right)=\sum_{z}\left[\Omega_{m}^{a_{z}}\delta_{c}^{2}\left(z\right)-\delta_{c\,\text{EdS}}^{2}\left(\frac{\delta_{{\rm v}}\left(z\right)}{\delta_{{\rm v}\,\text{EdS}}}\right)^{\beta}\right]^{2}\,.$ (3.11)
taking the sum in redshift from $z=0$ to $z=2$ with $10$ steps. We verified that concentrating the $z$ interval at lower $z$ does not change the results significantly. For $\Omega_{m}^{0}=0.3$, we get
@@
$D_{c}^{2}\left(z\right)\equiv\Omega_{m}^{a_{z}}\delta_{c}^{2}\left(z\right)\,\,\,\text{and}\,\,\,D_{{\rm v}}^{2}\left(z\right)\equiv\alpha\text{$\delta_{{\rm v}}^{\beta}\left(z\right)$}\,.$ (3.13)
In figure 2, we show the percent differences between $D_{{\rm v}}^{2}$ and $D_{c}^{2}$ for several values of $\Omega_{m}^{0}$. We observe that the difference increases for lower $\Omega_{m}^{0}$ values. This is expected because, when determining the parameter $\alpha$, we assumed EdS-limit, which is not well satisfied for low $\Omega_{m}^{0}$. On the other hand, the difference decreases for higher values of $\Omega_{m}^{0}$.
Figure 2: Difference between $D_{{\rm v}}^{2}$ and $D_{c}^{2}$ defined in (3.13) for several values of $\Omega_{m}^{0}$ indicated in the legend for $\Lambda$CDM model.
As can be seen, the representation of the redshift-dependent part of the effective threshold in the E-HMF is well described by a power-law function of $\delta_{{\rm v}}$, even for values of $\Omega_{m}^{0}$ distinct from the value used to determine the values of $\alpha$ and $\beta$. This indicates that, within $\Lambda$CDM, the virialization threshold can consistently capture the redshift-dependence described in the E-HMF. From the fact that $\Omega_{m}^{a_{z}}\delta_{c}^{2}\left(z\right)\propto\delta_{{\rm v}}^{4.37}$, we can expect that a more realistic SC model should provide a virialization threshold that grows faster at low-$z$, but this analysis is beyond the scope of this paper.

### v2

@@ Abstract
In a recent study by Euclid collaboration, the Halo Mass Function (HMF) has been fitted with accuracy better than $1\%$ for the $\Lambda$CDM model. Several parameters were introduced and fitted against N-body simulations, assuming the usual linearly extrapolated matter density contrast at the collapse time, $\delta_{c}$, as a basic threshold for halo formation. As a result, a new function that multiplies $\delta_{c}$ was introduced, producing an effective threshold that varies both with redshift and mass scale. We show that the redshift evolution of this effective threshold is similar to the one of the linear extrapolated matter density contrast at the virialization time, $\delta_{\rm v}$. Assuming the Euclid HMF as a fiducial model, we refit the Sheth-Tormen (ST) HMF using $\delta_{\rm v}$ as a threshold. This new fit improves the agreement between ST-HMF and the Euclid one with respect to Despali et al. (2016) fit, specially at high masses. Interestingly, the parameters $a$ and $p$ in this refit have values closer to the Press-Schechter limit of the ST-HMF, showing that the use of $\delta_{\rm v}$ can provide semi-analytical HMF less dependent on extra parameters, while providing better agreement with the Euclid HMF. Moreover, we analyze the consistency of the ST-HMF fitted with $\delta_{\rm v}$ in smooth dark energy models with time-varying equation of state, finding an overall good agreement with the evolution of halo abundances expected from the linear evolution of perturbations and the Euclid HMF extrapolated to these scenarios. These findings suggest that the use of $\delta_{\rm v}$ as a basic function to describe the threshold for halo formation can be a good guide when considering extrapolations for models beyond $\Lambda$CDM, which are typically harder to study in simulations. We also provide a fitting formula for $\delta_{\rm v}$ in the $\Lambda$CDM model and a code to compute it for smooth dark energy with equation of state described by the $w_{0}w_{a}$ parametrization.
@@ 1 Introduction
Since the first analytical proposal of a Halo Mass Function (HMF) by Press and Schechter (PS) [1], several semi-analytical and numerical studies on the topic have proposed alternative formulations, e.g., [2, 3, 4, 5, 6, 7, 8]. See [9] for an overview of several of these proposals. One of the key ingredients in the PS formulation, and in many others based on the Sheth-Tormen (ST) template [2], is the value of the critical density for the formation of halos, $\delta_{c}$, which mainly controls the exponential decay of the number of massive halos. Ideally, this parameter represents the value of the extrapolated linearly evolved matter contrast, $\delta_{m}$, above which halos are formed. For many years following the PS proposal, several distinct values for this threshold were used, e.g., [10, 11, 12]. Following the suggestions in [13, 14], the use of $\delta_{c}\simeq 1.686$, the value calculated for Einstein-de-Sitter (EdS) cosmology using the framework of the Spherical Collapse (SC) model [15, 16] at the moment of collapse, became standard.
@@
As suggested in [23, 24, 25], one can define another theoretical threshold for halo formation, $\delta_{{\rm v}}$, associated with virialization time. The definitions for the collapse and virialization threshold are:
@@
where $\delta_{m}^{L}$ is the linearly evolved matter density contrast at the redshift of collapse, $z_{c}$, or virialization, $z_{{\rm v}}$. Both of these quantities are computed in the framework of the SC model. For EdS, there is an analytical solution giving a constant $\delta_{{\rm v}}\simeq 1.583$. If one substitutes $\delta_{c}$ by $\delta_{{\rm v}}$ in HMFs, the threshold is reduced, as effectively done by the values of $a$ in the ST formulation, indicating that the virialization time can provide a better description of the abundance of halos, especially at the high-mass tail of HMFs.
@@
In this work, we will further explore the possibility of using $\delta_{{\rm v}}$ instead of $\delta_{c}$ in HMFs, focusing in the $\Lambda$CDM model. First, we will show that the redshift-dependent part of the E-HMF is in good agreement with a simple power-law function $\delta_{{\rm v}}(z)$, with precision better $0.7\%$ for $0.2\leq\Omega_{m0}\leq 0.4$, where $\Omega_{m0}=\Omega_{m}(z=0)$, indicating that the correspondence between the functions is not a coincidence for a particular value of $\Omega_{m0}$. Then we proceed with a reparametrization of the ST-HMF with the substitution $\delta_{c}\rightarrow\delta_{{\rm v}}$ and assuming the E-HMF as a fiducial model to fit its parameters, showing that this refit agrees better with E-HMF than the HMF fit given by [6]. As a result, we find that the refitted ST parameters $a$ and $p$ get closer to the values that reproduce the PS-HMF, indicating that the ST-HMF with $\delta_{{\rm v}}$ is less dependent on ad doc parameters. We also analyse how this new version of the ST-HMF behaves when considering smooth DE models with $w\neq-1$, showing it has a good overall agreement with the E-HMF. Consequently, one can more confidently use a correction function based on this new version of the ST-HMF to extrapolate the E-HMF to more complex cosmological models, provided that the corresponding modifications are made in the SC or virialization models used to compute $\delta_{\rm v}$.
The plan of this paper is as follows. In section 2 we review and discuss the collapse and virialization definitions for the thresholds. In section 3, we compare the redshift-dependent part of the E-HMF with $\delta_{{\rm v}}(z)$. In section 4, substituting $\delta_{c}(z)$ by $\delta_{{\rm v}}(z)$, we refit the ST-HMF, assuming the E-HMF as a fiducial model. In section 5, we analyse the consistency of the ST-HMF refitted with $\delta_{{\rm v}}(z)$ for smooth DE models with equation of state described by the $w_{0}w_{a}$ parametrization. We conclude in section 6 and provide a representation of $\delta_{{\rm v}}(z)$ in terms of a well-known fit for $\delta_{c}(z)$ for $\Lambda$CDM model in Appendix A, also indicating a repository where a code to compute $\delta_{{\rm v}}(z)$ assuming the $w_{0}w_{a}$ parametrization can be found.
Throughout this paper, we assume as baseline the $\Lambda$CDM model with parameters: $h=0.7$, $\Omega_{c0}h^{2}=0.125$, $\Omega_{b0}h^{2}=0.022$, $A_{s}=2\times 10^{-9}$, $n_{s}=0.965$, $\tau=0.06$, $m_{\nu}=0.06$ and $\Omega_{k0}=0$. In some examples, we vary $\Omega_{m0}$ by changing only $\Omega_{c0}h^{2}$. When analysing DE models described by the $w_{0}w_{a}$ parametrization, all the other parameters are kept fixed at the $\Lambda$CDM values. The computation of the matter power spectrum is made with CAMB [30].
@@ 2 Thresholds definitions
where $\rho_{m}\left(z\right)=\bar{\rho}_{m}\left(z\right)\left[1+\delta_{m}^{NL}\left(z\right)\right]$ is the total matter density. Note that $\Delta_{{\rm v}}$ is not the usual virialization overdensity, which is given by $\Delta=\rho_{m}\left(z_{\rm v}\right)/\bar{\rho}_{m}\left(z_{c}\right)\simeq 177.7\,.$ It is also worth mentioning that the spherical overdensity used to identify halos in the E-HMF analysis has a similar behaviour to the virialization overdensity defined in (2.5), see quantities $\Delta_{cc}$ and $\Delta_{v}$ in figure 1 of reference [28].
It is interesting to note that the virialization equation can be generalised to include other components, such as smooth or clustering DE [32, 33, 31, 26, 27], whereas the collapse definition is mostly insensitive to other ingredients that participate in the collapse process. Though in general $\delta_{m}^{NL}$ depends on the background evolution and other clustering components, near the collapse time, it is growing nearly exponentially, responding essentially to its own nonlinear value. Thus, the background evolution, and possibly other subdominant fluctuations, have very small impact on the value of $\delta_{c}(z)$, typically less than $1\%$ with respect to the EdS value [28].
@@
When DE begins to dominate the expansion, they become redshift-dependent. Whereas $\delta_{c}(z)$ has a slight decay when DE becomes important, $\delta_{{\rm v}}(z)$ presents a mild growth, see figure 1.
@@
$\delta_{cM}(z)=1.59+0.3114\times\ln\sigma_{8}\left(z\right)\,,$ (2.7)
@@
Assuming that the best representation of the effective threshold is given by $a(M,z)\delta_{c}^{2}(z)$, where $a(M,z)$ is given by (3.5), we clearly see that the usual $\delta_{c}(z)$ is both higher and has the opposite redshift variation. On the other hand, $\delta^{2}_{{\rm v}}(z)$ is closer to $a(M,z)\delta^{2}_{c}(z)$ and also increases at low-$z$. Note that the fit given in (2.7) was obtained in order to improve the accuracy of the halo model and the corresponding nonlinear matter power spectrum computed with it, and not necessarily the HMF, which is the quantity of interest in this work. But it is also clear that its behaviour is more compatible with $a(M,z)\delta_{c}^{2}(z)$ than $\delta_{c}^{2}(z)$.
In principle, $\delta_{cM}(z)$ also has good agreement with the effective threshold in the E-HMF. However, as can be seen in figure 1, the time variation of $\delta_{\rm v}(z)$ is more compatible with $a(M,z)\delta_{c}^{2}(z)$ than $\delta_{cM}(z)$. In particular, the $\ln\sigma_{8}\left(z\right)$ dependence in $\delta_{cM}(z)$ will not give a constant value at high-z, as should be in the EdS limit. Moreover, $\delta_{\rm v}(z)$ is described by a physical model, which can be improved and generalised for other cosmologies, whereas $\delta_{cM}(z)$ is based on a empirical fit. Note also that, in figure 1, the bare value of $\delta_{\rm v}(z)$ is shown, but $\delta_{cM}(z)$ has two adjusted parameters, so it is closer to the E-HMF effective threshold.
Figure 1: Evolution of $\delta_{c}^{2}(z)$, $\delta_{cM}^{2}(z)$, $\delta_{{\rm v}}^{2}(z)$ and $a(M,z)\delta_{c}^{2}(z)$ as a function of redshift in the $\Lambda$CDM model with $\Omega_{m0}=0.3$, $a_{z}=-0.0658$ and $a_{R}=0.7965$ (corresponding to a mass-scale $10^{14}M_{\odot}h^{-1}$ ).
@@ 3 The effective threshold in the E-HMF and $\delta_{{\rm v}}$
In this section, we analyse how the effective threshold of E-HMF is related to $\delta_{{\rm v}}(z)$ and define the main quantities of interest. The comoving number density of halos per mass interval is given by
@@
where $\nu f\left(\nu\right)$ is the multiplicity function, $\nu=\delta_{c}(z)/\sigma$, $\sigma^{2}$ is the smoothed variance of the matter power-spectrum, $P\left(k\right)$, given by
@@
$a(M,z)=a_{R}\Omega_{m}^{a_{z}}(z)=\left[a_{1}+a_{2}\left(\frac{d\ln\sigma}{d\ln R}+0.6125\right)^{2}\right]\Omega_{m}^{a_{z}}(z)\,,$ (3.5)
$p(M,z)=p_{1}+p_{2}\left(\frac{d\ln\sigma}{d\ln R}+0.5\right)\,,$ (3.6)
$q(M,z)=\left[q_{1}+q_{2}\left(\frac{d\ln\sigma}{d\ln R}+0.5\right)\right]\Omega_{m}^{q_{z}}(z)\,,$ (3.7)
@@
As can be seen, the function $a(M,z)$ always multiplies $\delta_{c}^{2}(z)$ and depends both on redshift and mass scale. The usual SC model is scale-independent, therefore it is not able to explain the variation with mass scale, which is encoded in $a_{R}$. Thus, we will focus on the redshift dependence of the effective threshold, $\Omega_{m}^{a_{z}}(z)\delta_{c}^{2}(z)$, and analyse its relation with $\delta_{{\rm v}}(z)$.
@@
The evolution of $a(M,z)\delta_{c}^{2}(z)$ is shown in figure 1, where we used $a_{R}=0.7965$, corresponding to a mass of $10^{14}M_{\odot}h^{-1}$. We can see that both $\delta_{{\rm v}}^{2}(z)$ and $a(M,z)\delta_{c}^{2}(z)$ increase at low-$z$, so they qualitatively represent the same trend for halo formation, i.e., a threshold that increases as $\Lambda$ becomes more important. Thus it is clear that the term $\Omega_{m}^{a_{z}}(z)$ inside $a(M,z)$ compensates the slight decay of $\delta_{c}(z)$ to better describe the abundance of simulated halos. Moreover, $\delta_{{\rm v}}^{2}(z)$ is closer to $a(M,z)\delta_{c}^{2}(z)$ then $\delta_{c}^{2}(z)$. Therefore, also from a direct analysis of the E-HMF, one can conclude that the physics associated with the definition of $\delta_{{\rm v}}(z)$ is more suitable than $\delta_{c}(z)$ because it is closer to the effective threshold and grows at low-$z$.
Relation between the effective threshold and $\delta_{\rm v}$
@@
Although $\delta_{\rm v}^{2}(z)$ and $a(M,z)\delta_{c}^{2}(z)$ have similar behaviours, it is also clear that there exists a mismatch between them. In figure 1, we can see that $a(M,z)\delta_{c}^{2}(z)$ grows faster at late times and has a lower value in the EdS limit. Therefore, it is important to check whether the similarity between them is not a coincidence for a specific set of cosmological parameters. Our goal now is to find a representation of the redshift-dependent part of the E-HMF effective threshold, given by $\Omega_{m}^{a_{z}}(z)\delta_{c}^{2}(z)$, in terms of $\delta_{{\rm v}}(z)$ and later check whether it is robust against variations of $\Omega_{m0}$. Let us analyse the following parametrization
$\Omega_{m}^{a_{z}}(z)\delta_{c}^{2}(z)=\alpha\delta_{{\rm v}}^{\beta}\left(z\right)\,,$ (3.8)
where $\alpha$ and $\beta$ are constants.
@@
It is important to note that the form $\alpha\delta_{{\rm v}}^{\beta}\left(z\right)$ will introduce a non-standard power-law dependence on the density threshold, i.e., distinct from power $2$. This is unusual in the context of statistical derivation of HMF [1, 13], because the squared filtered density contrast has different power than its variance. However, this reparametrization will not be implemented in the HMFs analysis in the next sections and will be used only to check the consistency between the functions when varying $\Omega_{m0}$.
Two other forms preserving the power two in $\delta_{\rm v}$ that could be used are: $\alpha\delta_{\rm v}^{2}(z)$ and $\alpha\Omega_{m}^{\beta}(z)\delta_{{\rm v}}^{2}\left(z\right)$. We checked that both of them are very precise when fitted for a fixed $\Omega_{m0}=0.3$, but the resulting fits are not robust when extrapolated to other values of $\Omega_{m0}$. Therefore, we will proceed with the expression in equation (3.8), which will bring an important insight about the virialization model.
@@
The constant $\alpha$ can be set by assuming that $\Omega_{m}^{a_{z}}(z)\delta_{c}^{2}(z)=\alpha\delta_{{\rm v}}^{\beta}(z)$ in the EdS limit, thus
@@
and $\beta$ is given by
$\beta=\ln\left(\frac{\Omega_{m}^{a_{z}}(z)\delta_{c}^{2}\left(z\right)}{\delta_{c\,\text{EdS}}^{2}}\right)/\ln\left(\frac{\delta_{{\rm v}}\left(z\right)}{\delta_{{\rm v}\text{\,EdS}}}\right)\,,$ (3.10)
which varies with $z$. For $\Omega_{m0}=0.3$ and $0\leq z\leq 2$, we have $4.514.05<\beta<$ 4.51.
@@
$\chi^{2}\left(\beta\right)=\sum_{z}\left[\Omega_{m}^{a_{z}}(z)\delta_{c}^{2}\left(z\right)-\delta_{c\,\text{EdS}}^{2}\left(\frac{\delta_{{\rm v}}\left(z\right)}{\delta_{{\rm v}\,\text{EdS}}}\right)^{\beta}\right]^{2}\,.$ (3.11)
taking the sum in redshift from $z=0$ to $z=2$ with $10$ steps. We verified that concentrating the $z$ interval at lower $z$ does not change the results significantly. For $\Omega_{m0}=0.3$, we get
@@
$D_{c}^{2}\left(z\right)\equiv\Omega_{m}^{a_{z}}(z)\delta_{c}^{2}\left(z\right)\,\,\,\text{and}\,\,\,D_{{\rm v}}^{2}\left(z\right)\equiv\alpha\text{$\delta_{{\rm v}}^{\beta}\left(z\right)$}\,,$ (3.13)
where $\alpha$ and $\beta$ are given by (3.12). In figure 2, we show the percent differences between $D_{{\rm v}}^{2}$ and $D_{c}^{2}$ for several values of $\Omega_{m0}$. We observe that the difference increases for lower $\Omega_{m0}$ values. This is expected because, when determining the parameter $\alpha$, we assumed EdS-limit, which is not well satisfied for low $\Omega_{m0}$. On the other hand, the difference decreases for higher values of $\Omega_{m0}$.
Figure 2: Difference between $D_{{\rm v}}^{2}$ and $D_{c}^{2}$ defined in (3.13) for several values of $\Omega_{m0}$ indicated in the legend for $\Lambda$CDM model.
As can be seen, the redshift-dependent part of the effective threshold in the E-HMF is well described by a power-law function of $\delta_{{\rm v}}$, even for values of $\Omega_{m0}$ distinct from the value used to determine the values of $\alpha$ and $\beta$. The fact that $\Omega_{m}^{a_{z}}(z)\delta_{c}^{2}\left(z\right)\propto\delta_{{\rm v}}^{4.37}(z)$ can be understood as manifestation of the non-universality of the E-HMF, introduced by the new fitting functions. Another possible interpretation is that a more realistic SC model should provide a virialization threshold that grows faster at low-$z$, in such a way that the new function could represent $D^{2}_{c}(z)$ with $\beta\simeq 2$, but this analysis is beyond the scope of this paper.
In order to get a HMF less dependent on empirical fitting function encoded in $D_{c}$, one could try to use the substitution $D_{c}\rightarrow D_{\rm v}$ in the E-HMF. As figure 2 indicates, this kind of reparametrization can be very accurate. We checked that it produces less than $5\%$ differences against the original E-HMF for $10^{13}M_{\odot}/h\leq M\leq 5\cdot 10^{15}M_{\odot}/h$, $0.25\leq\Omega_{m0}\leq 0.35$ and $0\leq z\leq 2$. However, we would be dealing with a reparametrization of a non-universal HMF, whose applicability to other cosmological models would be quite speculative. Therefore, in what follows, we will study how the use of $\delta_{\rm v}$ in the more universal ST-HMF can provide a better agreement with the E-HMF and later analyse how it can used to extrapolate the E-HMF to models beyond $\Lambda$CDM.